OpenStudy (anonymous):

3x + 5y = 7 2x - 3y = 11 Which method would work better for the solution? Linear combinations or substitution method?

OpenStudy (maddie82502):

so for example one method is called a Subtitution method and u have to use it first: Substitution method:Use the Substitution method to solve the system of equations. x + y = -4 x - y = 2 x + y = -4 x - y = 2 In the second equation, x= y+2 Putting that in the first.. (y+2) + y= -4 subtract two from each side, then divide each side by two to get y. so I think that is right but idk 4 sure